# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v1

@@ 3.1 PG–DPO as stochastic gradient ascent and BPTT–PMP correspondence
We write $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ for the (population) $q$-aggregated projection map. Theorem 3 below shows that, under a mild slab-wise local stability regime (i.e. a short-time contraction of $T$ on each $D_{k}$), small residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal deployable $\theta$-blind policy, up to discretization/Monte Carlo error. The proof combines a projection-map stability bound (Appendix C.1) with a slab-wise closure (Appendix C.2), in the same spirit as the slab analyses used in huh2025breaking.
@@
Assume the uniform invertibility/stability conditions of Proposition 1 (Appendix C.1) hold on $D$ for the relevant aggregated curvature terms and for the estimator perturbations constructed under $\pi^{\mathrm{warm}}$.
@@
and assume the slab-wise Lipschitz gain of Appendix C.2 holds: there exist constants $\bar{L}_{A},\bar{L}_{G}>0$ such that for every slab $D_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
@@
Let $\widehat{\pi}^{\mathrm{agg,mix}}$ be the mixed-moment projected policy (67) computed from BPTT/Monte Carlo estimates under $\pi^{\mathrm{warm}}$, and let $\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}$ be the population residual (68). Then there exists $C_{2}>0$ such that
$\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)}\;\leq\;\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}\;+\;C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}).$ (71)
Moreover, under the perturbative regime of Proposition 1, one may take for example $C_{2}:=2\kappa+4\kappa^{2}M_{G}$.
@@
See Appendix C.3. ∎
3.3 Coupling stage 1 and stage 2: residual projection and interactive distillation
We keep the ex–ante objective (56) and the $\theta$-blind deployability constraint throughout. Stage 2 is not a separate optimization problem: it reuses the current stage 1 policy as a warm-up control under which the (costate-based) projection ingredients are estimated, and then applies a $q$-aggregated Pontryagin projection as a post-processing map.
This subsection records two couplings between the two stages, each with a distinct role. First, we implement the projected rule in a residual (control-variate) form, which is algebraically equivalent to the direct projection but typically reduces Monte Carlo variance and improves numerical stability in high dimensions. Second, we use the projected output as a teacher signal via interactive distillation. Beyond acting as an optimization aid, distillation serves an amortization purpose: stage 2 projection can be accurate but Monte-Carlo intensive, whereas a distilled student policy can approximate the projected rule with a single forward pass at stage 1 inference cost.
3.3.1 Control-variate (residual) form of the projected rule
Recall the mixed-moment projected rule (67). In high dimensions, Monte Carlo noise in the projection inputs can be non-negligible, and solving a linear system with $\widehat{A}_{t}$ can amplify this noise. A convenient stabilization is to compute the same projected rule in a residual form around the warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$.
and the aggregated residual (the quantity we actually solve against)
Whenever $\widehat{A}_{t}(t,x,y)$ is invertible, the projected rule admits the identity
which is an algebraic rewriting of (67) (hence it does not change the target). Its practical value is variance reduction: when the warm-up policy is already close to a projected fixed point on the working domain, the residual $\widehat{r}_{\mathrm{FOC}}$ tends to be small, and it often concentrates faster because the ingredients entering $\widehat{A}_{t}\pi_{\varphi^{\mathrm{warm}}}$ and $\widehat{G}_{t}^{\mathrm{mix}}$ are computed from the same Monte Carlo pool and partially cancel.
3.3.2 Interactive distillation: projection-guided training and amortized deployment
Let $\pi_{\varphi}$ be the trainable stage 1 policy network. At intermittent refresh times, we freeze a lagged copy $\pi_{\varphi^{-}}$ and run stage 2 under $\pi_{\varphi^{-}}$ to construct a $q$-aggregated projected teacher. This coupling serves two purposes. During training it provides projection-guided targets that can stabilize and accelerate stage 1 optimization; after training it amortizes the expensive projection by distilling it into a fast deployable policy network.
In residual form (74), the teacher is the $\theta$-blind map
where $\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}:=\widehat{A}_{t}^{\varphi^{-}}\pi_{\varphi^{-}}+\widehat{G}_{t}^{\mathrm{mix},\varphi^{-}}$ is computed using the mixed-moment $q$-aggregation under the lagged policy. We then train $\pi_{\varphi}$ by combining the original ex–ante objective with a proximity term to this teacher on the working domain:
where $\mu$ is the working-domain sampling measure and $\lambda\geq 0$ controls the strength of projection guidance. The operator $\mathrm{stopgrad}(\cdot)$ indicates that gradients are not propagated through stage 2; once computed from $\pi_{\varphi^{-}}$, the teacher is treated as fixed.
In practice, $\varphi^{-}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}$ are refreshed on a slower timescale than the stage 1 gradient steps: we hold $\varphi^{-}$ fixed for several updates of $\varphi$ under (76), then set $\varphi^{-}\leftarrow\varphi$ and recompute the teacher. A practical schedule is to start with $\lambda=0$ (pure PG–DPO) and increase $\lambda$ only after basic projection checks on the working domain (e.g., residual magnitudes and curvature/denominator stability) indicate that the stage 2 map has become reliable. Moreover, to avoid injecting noisy teacher targets early in training or on pathological regions of the domain, we may apply projection guidance only on states where the projection checks certify reliability (an “adaptive teacher selection”); implementation details are deferred to the appendix (Appendix D).
@@ 4 Breaking the Dimensional Barrier under Drift Uncertainty
This section instantiates the decision-time static Gaussian drift-uncertainty benchmark in Section 2.3.1 and uses its closed-form constant-portfolio $q$-reference as an analytic target. Nature draws a fixed latent drift $\theta\sim q$ at $t=0$ and keeps it constant over $[0,T]$, while the investor cannot observe $\theta$ and must deploy a single $\theta$-blind policy under an ex–ante CRRA objective. Because the benchmark admits a transparent decision-time reference, we can measure accuracy directly via decision-time RMSE, rather than relying only on realized utility.
Our goal is to test whether Pontryagin-guided learning and projection remain stable as the number of assets grows. We generate APT-style covariance structures and sweep dimensions $d\in\{5,10,50,100\}$ under both aligned uncertainty ($P=s\Sigma$) and a misaligned geometry that rotates uncertainty away from market risk directions. We compare Stage 1 (PG–DPO) to Stage 2 (Pontryagin projection) (and, when applicable, amortized variants via interactive distillation) under matched simulation budgets that scale linearly with $d$.
@@ 4.1 Benchmark market and evaluation protocol
$\theta$-blind deployability (and what uses $\theta$). Throughout Section 4, all reported policies are deployable and $\theta$-blind: the control is a function of observable state only (here, decision-time evaluation uses $t=0$ and $X_{0}$), and never takes the realized latent premium $\theta$ as an input. The latent $\theta\sim q$ is sampled only inside the simulator to generate trajectories and to form Monte Carlo averages that approximate $q$-expectations (notably in Stage 2 projection). Any $\theta$-indexed objects (when referenced elsewhere) are used only for offline diagnostics and are not part of the deployable decision rule.
@@
APT-style factor construction of $(m,\Sigma)$. We construct the mean premium and covariance via a low-dimensional factor representation. Let $W^{f}$ be a $k_{\Sigma}$-dimensional Brownian motion (factor shocks) and $W^{\varepsilon}$ a $d$-dimensional Brownian motion (idiosyncratic shocks), independent of $W^{f}$. We write excess returns as
@@
where $\widetilde{B}$ is generated independently of $B$ (or explicitly orthogonalized against the span of $B$ to enforce large principal angles). The factor term is rescaled so that its overall magnitude matches the aligned case under the same $s$ (e.g., by matching $\mathrm{tr}(P)$ or $\|P\|_{F}$ up to the shared diagonal component). This geometry increases heterogeneity across $\theta\sim q$ and makes mixed-moment estimation and subsequent linear-algebra steps more fragile, providing a stringent scalability test.
Experiment grid and simulation budgets. We vary the number of assets over $d\in\{5,10,50,100\}$ and sweep three uncertainty magnitudes $s\in\{10^{-3},10^{-2},10^{-1}\}$, for both aligned and misaligned geometries. To keep Monte Carlo noise comparable across dimensions, we use linear-in-$d$ sampling budgets: a base regime with $N_{\mathrm{MC}}=100\cdot d$ paths and a high regime with $N_{\mathrm{MC}}=400\cdot d$ (where $N_{\mathrm{MC}}$ denotes the per-update or per-estimator path budget, depending on the stage). All methods share the same discretization scheme (Euler) and action constraints; implementation details (network architecture, optimizer settings, and exact sampling conventions for Stage 1 vs. Stage 2) are reported in the implementation appendix and code release.
Analytic reference and decision-time evaluation. In the static Gaussian benchmark, the analytic decision-time reference under constant portfolios is available in closed form. We use this closed-form rule only as an external decision-time target for evaluation; training does not impose the constant-portfolio restriction, and all methods learn from simulated trajectories over $[0,T]$ under the same $\theta$-blind constraint. For $\gamma>1$ we use the CRRA reference (38) and, for $\gamma=1$, the log-utility reference (31). We evaluate each method at $t=0$ on a fixed grid $\{(X_{0}^{(i)},T^{(i)})\}_{i=1}^{N_{\mathrm{eval}}}$ and report RMSE to the analytic reference:
$\mathrm{RMSE}(u_{0},\pi_{q,\gamma}^{\mathrm{const}}):=\bigg(\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\big\|u_{0}(X_{0}^{(i)},T^{(i)})-\pi_{q,\gamma}^{\mathrm{const}}(T^{(i)})\big\|^{2}\bigg)^{1/2},$ (82)
where $u_{0}(\cdot)$ denotes the decision-time action prescribed by the method (deployable $\theta$-blind output). With the benchmark fixed and with $(m,\Sigma,P)$ constructed as in (79)–(81), the remaining subsections compare baseline Stage 1 PG–DPO, post-hoc Stage 2 P–PGDPO projection, and interactive distillation under matched simulation budgets.
@@ 4.2 High-dimensional CRRA benchmark: projection and amortization
Figure 1: Decision-time RMSE at $t=0$ versus dimension $d$ (log scale), summarized by a tail median over the late-training window (computed from the last evaluation snapshots). Rows: uncertainty magnitude $s\in\{10^{-3},10^{-2},10^{-1}\}$. Columns: aligned vs. misaligned geometry. Curves compare Stage 1 (deployable) and Stage 2 (post-hoc projection), with and without interactive distillation. Solid vs. dashed lines correspond to MC base ($100\cdot d$) vs. high ($400\cdot d$) budgets.
Protocol and summary statistic. We consider the CRRA benchmark with $\gamma=2$ under Gaussian drift uncertainty $q$ and evaluate against the analytic constant $q$-reference (38). We track (i) the Monte-Carlo objective estimate $\widehat{J}$ during training and (ii) the decision-time error at $t=0$ via RMSE (82). Because stochastic optimization produces non-monotone and noisy RMSE curves, we summarize each condition by a robust tail median: the median RMSE over the final evaluation snapshots in the late-training window. Unless stated otherwise, the projection/teacher direction uses the mixed-moment ($p_{\theta}$) aggregation.
What is compared in Figure 1. Stage 1 (PG–DPO; Section 3.1) trains a deployable $\theta$-blind policy $\pi_{\varphi}$ by maximizing $\widehat{J}$ via pathwise gradients. Stage 2 (P–PGDPO; Section 3.2) applies a $q$-aggregated Pontryagin projection to a Stage 1 checkpoint; we use the residual form of Section 3.3.1. Interactive distillation (Section 3.3.2) treats the Stage 2 projected control as a teacher signal and amortizes it back into a deployable Stage 1 policy network.
@@
Stage 2 projection versus amortization: scaling with dimension. Aligned geometry. For small and moderate uncertainty ($s=10^{-3},10^{-2}$), Stage 2 delivers a sharp reduction in decision-time error across all tested dimensions, bringing RMSE down to the $10^{-5}$$10^{-4}$ range, while Stage 1 policies remain around $10^{-3}$. Interactive distillation consistently improves the deployable policy (Stage 1 (distill.) below Stage 1) while leaving Stage 2 essentially unchanged, confirming the intended division of labor: Stage 2 supplies a structured stationarity-correction signal, and distillation reduces the policy-class approximation/optimization gap by injecting that signal into $\pi_{\varphi}$.
@@
Misaligned geometry. The picture becomes more heterogeneous. For small to moderate uncertainty ($s=10^{-3},10^{-2}$), Stage 2 still improves decision-time RMSE at small $d$, but its advantage shrinks with dimension and can approach the $10^{-3}$ level by $d=100$. For the largest uncertainty scale ($s=10^{-1}$), Stage 1 becomes markedly less reliable, whereas Stage 2 remains substantially better, indicating that projection can act as a stabilizing correction even when end-to-end learning is stressed. Across settings, the base and high MC budgets tend to yield similar tail-median RMSE curves, suggesting that linear-in-$d$ scaling of simulation budgets is sufficient for stable comparisons in this benchmark.
@@
Mechanism: why misalignment can reduce projection gains. To explain when and why the projection gains shrink, we analyze Stage 2 diagnostic statistics reported in Appendix E; see Figures 36. The diagnostics indicate that the degradation under misalignment is driven primarily by increased stationarity residuals and curvature mismatch, rather than by catastrophic denominator sign failures: (i) the Stage 2 residual norm grows with dimension and becomes especially large in the hardest misaligned regime, (ii) the projection denominator magnitude stays away from zero at typical quantiles, and (iii) the bad-sign fraction remains negligible, while (iv) the effective curvature statistic $\kappa$ stays near the nominal $1/\gamma$ reference in easy regimes but can deviate substantially in the hardest misaligned/high-uncertainty setting. These patterns are consistent with the geometric explanation: when $P$ and $\Sigma$ do not commute, the inverse operations implicit in projection mix directions and can amplify Monte-Carlo errors in mixed-moment quantities (e.g., $\mathbb{E}[p_{1}\theta]$), especially as $d$ increases.
Figure 2: Pathwise sanity check at $d=100$ under common random numbers (same sampled $\theta$ and Brownian increments). Top: aligned geometry. Bottom: misaligned geometry. Each panel shows $\log X_{t}$ trajectories induced by the warm Stage 1 policy (PGDPO), the online Stage 2 P–PGDPO teacher (residual form), and the analytic $q$-reference.
@@
Pathwise sanity check. Figure 2 complements the decision-time RMSE with a trajectory-level view under common random numbers. In the aligned case, the online Stage 2 teacher tracks the analytic $q$-reference closely along a realized path and reduces the deviation $\Delta\log X_{t}$ relative to the warm Stage 1 policy. In the misaligned case, the teacher can deviate more noticeably under the same common-noise protocol, mirroring the reduced projection advantage in the hardest regimes of Figure 1 and motivating amortization/reliability mechanisms in interactive distillation.
Overall, the benchmark highlights a separation of roles. Stage 2 projection supplies a structured stationarity-correction signal that is particularly effective under aligned uncertainty, and interactive distillation amortizes this signal into a fast deployable Stage 1 policy. Under misalignment, projection can become more sensitive as $d$ and $s$ grow, consistent with diagnostic evidence of increased residuals and curvature mismatch; nevertheless, amortization remains a robust route to improving deployable policies under fixed simulation budgets.

### v2

@@ 3.1 PG–DPO as stochastic gradient ascent and a conditional BPTT–PMP correspondence
We write $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ for the (population) $q$-aggregated projection map. Theorem 3 below shows that, under a mild slab-wise local stability regime (i.e. a short-time contraction of $T$ on each $D_{k}$), small residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal deployable $\theta$-blind policy, up to discretization/Monte Carlo error. The proof combines a projection-map stability bound (Appendix D.1) with a slab-wise closure (Appendix D.2), in the same spirit as the slab analyses used in Huh et al. (2025b).
Stage 1: PG–DPO
Sample $\theta\sim q$ in the simulator; Euler rollout; BPTT; update $\varphi$.
Output: $\varphi^{\mathrm{warm}}$.
Stage 2: $q$-aggregated projection
Under $\pi_{\varphi^{\mathrm{warm}}}$, estimate adjoint blocks; aggregate over $\theta$; solve $A\pi=-G$.
Output: $\widehat{\pi}^{\mathrm{agg,mix}}$.
Optional couplings
Residual-form projection; interactive distillation (teacher refresh).
Figure 1: Two-stage pipeline of Section 3. A detailed schematic (including the coupling mechanisms and implementation notes) is provided in Appendix F, Fig. LABEL:fig:sec3-pipeline-sideways.
@@
Assume the uniform invertibility/stability conditions of Proposition 1 (Appendix D.1) hold on $D$ for the relevant aggregated curvature terms and for the estimator perturbations constructed under $\pi^{\mathrm{warm}}$ (in particular, that $\widehat{A}_{t}$ remains a small perturbation of $A_{\pi^{\mathrm{warm}}}$ on $D$ under the adopted stabilizers).
@@
and assume the slab-wise Lipschitz gain of Appendix D.2 holds: there exist constants $\bar{L}_{A},\bar{L}_{G}>0$ such that for every slab $D_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
@@
Let $\widehat{\pi}^{\mathrm{agg,mix}}$ be the mixed-moment projected policy (59) computed from BPTT/Monte Carlo estimates under $\pi^{\mathrm{warm}}$ (with stabilized inversion as needed), and let $\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}$ be the population residual (60). Then there exists $C_{2}>0$ such that
$\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)}\;\leq\;\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}\;+\;C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$ (65)
where one admissible explicit choice of $\delta_{\mathrm{BPTT}}$ is given in (62). Moreover, under the perturbative regime of Proposition 1, one may take for example $C_{2}:=2\kappa+4\kappa^{2}M_{G}$.
@@
See Appendix D.3. ∎
Coupling and amortization.
Figure 1 summarizes the overall two-stage design. Stage 2 is used purely as a warm-started post-processing map: it reuses the current stage 1 policy, estimates the adjoint blocks under frozen $\theta\sim q$, and outputs a single deployable $\theta$-blind projected rule. For numerical stability and amortized deployment, we additionally consider (i) a residual/control-variate form of the projection and (ii) interactive distillation of the projected rule into the policy network. These practical coupling mechanisms are described in Appendix F.
@@ 4 Breaking the Dimensional Barrier under Drift Uncertainty
This section instantiates the decision-time static Gaussian drift-uncertainty benchmark from Section 2.3.1 and uses its closed-form constant-portfolio $q$-reference as an analytic target. Nature draws a latent drift $\theta\sim q$ at $t=0$ and keeps it constant over $[0,T]$, while the investor cannot observe $\theta$ and must deploy a single $\theta$-blind policy under an ex–ante CRRA objective. Because the benchmark admits a transparent decision-time reference, we can measure accuracy directly via a decision-time Euclidean RMSE to the analytic reference, rather than relying only on realized utility.
Our goal is to test whether Pontryagin-guided learning and projection remain stable as the number of assets grows. We generate APT-style covariance structures and sweep dimensions $d\in\{5,10,50,100\}$ under both aligned uncertainty ($P=s\Sigma$) and a misaligned geometry that rotates uncertainty away from market risk directions. We compare Stage 1 (PG–DPO) to Stage 2 (Pontryagin projection), and (when applicable) amortized variants via interactive distillation, under Monte Carlo budgets that scale linearly with $d$.
@@ 4.1 Benchmark market and evaluation protocol
$\theta$-blind deployability (and what uses $\theta$). Throughout Section 4, all reported policies are deployable and $\theta$-blind: the control is a function of observable state only (here, decision-time evaluation uses $(t,X_{0})$ with fixed horizon $T$), and never takes the realized latent premium $\theta$ as an input. The latent $\theta\sim q$ is sampled only inside the simulator to generate trajectories and to form Monte Carlo averages that approximate $q$-expectations (notably in Stage 2 projection). Any $\theta$-indexed objects (when referenced elsewhere) are used only for offline diagnostics and are not part of the deployable decision rule.
@@
APT-style factor construction of $(m,\Sigma)$. We construct the mean premium and covariance via a low-dimensional factor representation in the spirit of arbitrage pricing theory (APT) Ross (1976). Let $W^{f}$ be a $k_{\Sigma}$-dimensional Brownian motion (factor shocks) and $W^{\varepsilon}$ a $d$-dimensional Brownian motion (idiosyncratic shocks), independent of $W^{f}$. We write excess returns as
@@
where $\widetilde{B}$ is generated independently of $B$ (and can be orthogonalized against the span of $B$ to enforce large principal angles). The factor term is rescaled to match the aligned factor magnitude under the same $s$; concretely, writing $\Sigma_{\mathrm{fac}}:=B\Sigma_{f}B^{\top}$ and $\widetilde{\Sigma}_{\mathrm{fac}}:=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}$, we rescale $\widetilde{\Sigma}_{\mathrm{fac}}$ so that $\mathrm{tr}(\widetilde{\Sigma}_{\mathrm{fac}})=\mathrm{tr}(s\,\Sigma_{\mathrm{fac}})$ while keeping the diagonal term fixed as $s\,\mathrm{diag}(D)$. This geometry increases heterogeneity across $\theta\sim q$ and makes mixed-moment estimation and subsequent linear-algebra steps more fragile, providing a stringent scalability test.
Experiment grid and simulation budgets. We vary the number of assets over $d\in\{5,10,50,100\}$ and sweep three uncertainty magnitudes $s\in\{10^{-3},10^{-2},10^{-1}\}$ for both aligned and misaligned geometries. To keep Monte Carlo noise comparable across dimensions, we use linear-in-$d$ trajectory budgets: a base regime with $N_{\mathrm{MC}}=100\cdot d$ rollouts and a high regime with $N_{\mathrm{MC}}=400\cdot d$ rollouts. Here $N_{\mathrm{MC}}$ denotes the number of independent simulated trajectories used to approximate the relevant $q$-expectations (in Stage 1 for stochastic gradient estimates, and in Stage 2 for costate/mixed-moment estimators at queried states). Each trajectory draws its own latent $\theta\sim q$ at time 0 and its own Brownian path thereafter; we do not use a separate inner loop of $\theta$-samples per trajectory. All methods share the same discretization scheme (Euler) and the same market instances within each $d$; implementation details (network architecture, optimizer settings, and exact sampling conventions for Stage 1 vs. Stage 2) are reported in the implementation appendix and code release.
Analytic reference and decision-time evaluation. In the static Gaussian benchmark, the analytic decision-time reference under constant portfolios is available in closed form. We use this closed-form rule only as an external decision-time target for evaluation; training does not impose the constant-portfolio restriction, and all methods learn from simulated trajectories over $[0,T]$ under the same $\theta$-blind constraint. For $\gamma>1$ we use the CRRA reference (31). We evaluate each method at $t=0$ with a fixed horizon $T=1.5$ on a fixed grid $\{X_{0}^{(i)}\}_{i=1}^{N_{\mathrm{eval}}}$ and report Euclidean RMSE to the analytic reference:
$\mathrm{RMSE}(u_{0},\pi_{q,\gamma}^{\mathrm{const}}):=\bigg(\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\big\|u_{0}(X_{0}^{(i)},T)-\pi_{q,\gamma}^{\mathrm{const}}(T)\big\|^{2}\bigg)^{1/2},$ (71)
where $\|\cdot\|$ is the Euclidean norm on $\mathbb{R}^{d}$ and $u_{0}(\cdot)$ denotes the decision-time action prescribed by the method (deployable $\theta$-blind output). With the benchmark fixed and with $(m,\Sigma,P)$ constructed as in (68)–(70), the remaining subsections compare baseline Stage 1 PG–DPO, post-hoc Stage 2 P–PGDPO projection, and interactive distillation under matched simulation budgets.
@@ 4.2 High-dimensional CRRA benchmark: projection and amortization
Figure 2: Decision-time Euclidean RMSE at $t=0$ versus dimension $d$ (log scale), summarized by a tail median over the late-training window. Concretely, for each condition we take the median RMSE over the last evaluation snapshots recorded during training. Rows: uncertainty magnitude $s\in\{10^{-3},10^{-2},10^{-1}\}$. Columns: aligned vs. misaligned geometry. Curves compare Stage 1 (deployable) and Stage 2 (post-hoc projection), with and without interactive distillation. Solid vs. dashed lines correspond to MC base ($100\cdot d$) vs. high ($400\cdot d$) trajectory budgets.
Protocol and summary statistic.
We consider the CRRA benchmark with $\gamma=2$ under Gaussian drift uncertainty $q$ and evaluate against the analytic constant $q$-reference (31). We track (i) the Monte-Carlo objective estimate $\widehat{J}$ during training and (ii) the decision-time error at $t=0$ via RMSE (71) with fixed horizon $T=1.5$. Because stochastic optimization produces non-monotone and noisy RMSE curves, we summarize each condition by a robust tail median: the median RMSE over the final evaluation snapshots in the late-training window. Unless stated otherwise, the projection/teacher direction uses the $p_{\theta}$ aggregation (Section 3.2).
What is compared in Figure 2. Stage 1 (PG–DPO; Section 3.1) trains a deployable $\theta$-blind policy $\pi_{\varphi}$ by maximizing $\widehat{J}$ via pathwise gradients. Stage 2 (P–PGDPO; Section 3.2) applies a $q$-aggregated Pontryagin projection to a Stage 1 checkpoint; we use the residual form of Section F.1. Interactive distillation (Section F.2) treats the Stage 2 projected control as a teacher signal and amortizes it back into a deployable Stage 1 policy network.
@@
Figure 3: Pathwise sanity check at $d=100$ under common random numbers. Black: analytic Gaussian constant-fraction benchmark evaluated with the remaining horizon $\tau=T-t$ (fixed $q$). Blue: Stage 1 PG–DPO. Orange: on-the-fly Stage 2 P–PGDPO teacher (computed via Monte Carlo projection at visited states). Top/bottom: aligned/misaligned.
@@
Stage 2 projection versus amortization: scaling with dimension. For the aligned geometry, small and moderate uncertainty ($s=10^{-3},10^{-2}$) yield a sharp Stage 2 reduction in decision-time error across all tested dimensions, bringing RMSE down to the $10^{-5}$$10^{-4}$ range, while Stage 1 policies remain around $10^{-3}$. Interactive distillation consistently improves the deployable policy (Stage 1 (distill.) below Stage 1) while leaving Stage 2 essentially unchanged, confirming the intended division of labor: Stage 2 supplies a structured stationarity-correction signal, and distillation reduces the policy-class approximation/optimization gap by injecting that signal into $\pi_{\varphi}$.
@@
In contrast, under the misaligned geometry the picture becomes more heterogeneous. For small to moderate uncertainty ($s=10^{-3},10^{-2}$), Stage 2 still improves decision-time RMSE at small $d$, but its advantage shrinks with dimension and can approach the $10^{-3}$ level by $d=100$. For the largest uncertainty scale ($s=10^{-1}$), Stage 1 becomes markedly less reliable, whereas Stage 2 remains substantially better, indicating that projection can act as a stabilizing correction even when end-to-end learning is stressed. Across settings, the base and high MC budgets tend to yield similar tail-median RMSE curves, suggesting that linear-in-$d$ scaling of simulation budgets is sufficient for stable comparisons in this benchmark.
@@
Mechanism: why misalignment can reduce projection gains. To explain when and why the projection gains shrink, we analyze Stage 2 diagnostic statistics reported in Appendix G; see Figures 47. The diagnostics indicate that the degradation under misalignment is driven primarily by increased stationarity residuals and curvature mismatch, rather than by catastrophic denominator sign failures: (i) the Stage 2 residual norm grows with dimension and becomes especially large in the hardest misaligned regime, (ii) the projection denominator magnitude stays away from zero at typical quantiles, and (iii) the bad-sign fraction remains negligible, while (iv) the effective curvature statistic $\kappa$ stays near the nominal $1/\gamma$ reference in easy regimes but can deviate substantially in the hardest misaligned/high-uncertainty setting. These patterns are consistent with the geometric explanation: when $P$ and $\Sigma$ do not commute, the inverse operations implicit in projection mix uncertainty directions that are poorly aligned with risk directions, which can amplify estimation noise in the projected correction as $d$ grows.
@@
Pathwise sanity check. Figure 3 complements the decision-time RMSE with a trajectory-level view under common random numbers. In the aligned case, the on-the-fly Stage 2 teacher tracks the analytic $q$-reference closely along a realized path and reduces the deviation $\Delta\log X_{t}$ relative to the warm Stage 1 policy. In the misaligned case, the teacher can deviate more noticeably under the same common-noise protocol, mirroring the reduced projection advantage in the hardest regimes of Figure 2 and motivating amortization/reliability mechanisms in interactive distillation.